# Past Movements-Guided Motion Representation Learning for Human Motion Prediction

### v1

@@ Abstract
Human motion prediction based on 3D skeleton is a significant challenge in computer vision, primarily focusing on the effective representation of motion. In this paper, we propose a self-supervised learning framework designed to enhance motion representation. This framework consists of two stages: first, the network is pretrained through the self-reconstruction of past sequences, and the guided reconstruction of future sequences based on past movements. We design a velocity-based mask strategy to focus on the joints with large-scale moving. Subsequently, the pretrained network undergoes finetuning for specific tasks. Self-reconstruction, guided by patterns of past motion, substantially improves the model’s ability to represent the spatiotemporal relationships among joints but also captures the latent relationships between past and future sequences. This capability is crucial for motion prediction tasks that solely depend on historical motion data. By employing this straightforward yet effective training paradigm, our method outperforms existing state-of-the-art methods, reducing the average prediction errors by 8.8% across Human3.6M, 3DPW, and AMASS datasets. The code is available at https://github.com/JunyuShi02/PMG-MRL.
@@ 1 Introduction
3D skeleton-based motion prediction stands as a fundamental task in computer vision, with the objectives of generating consecutive and unseen future motion sequences based on past observations. This technique finds widespread application in domains such as human-computer interaction, autonomous driving, and motion analysis. The key to achieve accurate predictions of future motions lies in effectively modeling motion sequences, particularly the spatiotemporal representation of moving joints. However, due to the inherent properties of human kinematics, joints exhibit complex hierarchical rotational relationships. The motion trajectories of human joints also display a certain degree of randomness. Consequently, effectively capturing the spatiotemporal relationships between joints has perennially posed a challenging problem.
In the early stages, researchers utilized probabilistic models, such as Gaussian processes[38] and nonlinear Markov[19] models, to predict future motion changes based on historical motion sequences. However, these methods are computationally intensive and make strong assumptions about data distribution, which leads to models that are inadequate for real-world prediction scenarios. With the emergence of large-scale datasets[17, 37, 27] and advancements in hardware computing performance, motion prediction tasks started employing deep learning methods based on Recurrent Neural Networks (RNNs)[11, 16, 9, 14, 31]. RNN-based models effectively capture the dependencies between frames in a sequence, but the step-by-step generation approach results in the accumulation of prediction errors. The rise of Graph Neural Networks (GNNs) has brought feed-forward prediction paradigms into the motion prediction[26, 3, 10, 30, 34]. GNNs are well-suited for handling the non-Euclidean distribution properties of joints in spatial dimensions, enabling effective modeling of motion’s spatial information. Moreover, methods based on Transformers and Multi Layer Perceptrons (MLPs) have also achieved success in motion prediction tasks[1, 12, 2].
Overall, the mainstream paradigms for 3D skeleton-based motion prediction can be categorized into two main categories: 1) extracting features from historical motion sequences using an encoder and then generating predicted motions by a decoder[20, 11, 23] (as shown in Fig. 1(a)); 2) directly generating predicted motions using a feed-forward network[28, 26, 22] (as shown in Fig. 1(b)). However, the potential of self-supervised learning paradigms based on pretraining has not been sufficiently explored and applied in the field of motion prediction. This limits the ability of models to fully learn the spatiotemporal representations of motion sequences. AuxFormer[41] improves prediction performance by incorporating denoising tasks and masked feature prediction as auxiliary tasks during training (as shown in Fig. 1(c)). Nevertheless, parallel multi-task prediction encounters challenges such as interference with the main task (future motion prediction) and difficulties in balancing loss weights. Moreover, conducting masked feature prediction directly on the entire sequence lacks an exploration of the relationship between historical and future sequences.
In this paper, we propose a novel motion representation learning method for motion prediction tasks, which comprises two stages: pretraining and finetuning. In the pretraining stage, we apply masking to the joints of future motion sequences and utilize the complete past sequence features to guide the reconstruction of future sequences. This approach not only enhances the model’s reconstruction performance through past sequence features but also allows the model to learn the underlying relationship between historical and future sequences during the pretraining stage. Simultaneously, we also conduct self-reconstruction training based on masking on the past sequences. When applying the mask, we mask the joints with the largest motion changes between adjacent frames (i.e., the set of joints with the highest velocity). These joints serve as the central representation of the primary motion changes. Compared to random masking, this approach avoids learning motion patterns of joints with lower information content. For instance, in the ”Eating” action, motion changes are concentrated in the upper limb part, implying that we should focus on the representation learning of the upper limb part (including hand, head, and other joints), rather than the lower limb part with smaller changes.
Motion data contains both temporal and spatial dimensions. Given the outstanding performance of Transformers in temporal modeling tasks and spatial modeling tasks, as well as prior work on using Transformers for motion prediction tasks [41, 32, 1], we employ spatiotemporal Transformers to model motion sequences. Specifically, our method includes two components based on Transformers: the Past Motion Encoder (PME) and the Future Motion Predictor (FMP). The former comprises SpatioTemporal Attention (ST-Attention) modules, while the latter consists of ST-Attention modules and Past Motion Guided Attention (PMG-Attention) modules that fuse historical and future information. By pretraining the parameters of the PME and FMP, and finetuning them for the prediction task, our method significantly outperforms other state-of-the-art methods in both long-term and short-term predictions using a straightforward model architecture. It is worth noting that we observe a correlation between the performance improvement of our method and the increase in dataset size, which aligns with the characteristic of self-reconstruction learning requiring large-scale data.
The key contributions of our work can be summarized as follows:
Figure 1: (a) Prediction based on Encoder-Decoder paradigm[20] [23]. (b) Prediction based on Feed-forward paradigm[24] [28]. (c) Enhancing prediction based on auxiliary task[6] [41]. (d) Our Pretrain-Finetune prediction paradigm.
  • We propose a human motion representation learning method based on self-reconstruction pretraining paradigm for motion prediction tasks. This method helps the model learn the intrinsic spatiotemporal structure and features of motion sequences.
  • We use past motion features to guide the self-reconstruction of future sequences, which not only improves self-reconstruction performance but also helps the model understand the underlying relationship between past and future sequences in the motion prediction tasks.
@@
  • We conduct extensive experiments and demonstrate that our method reduces the average prediction errors by 8.8% compared to state-of-the-art methods on the 3DPW, Human3.6M, and AMASS datasets.
@@ 2 Related Work
RNN-based Methods
Early motion prediction methods employed probability model-based approaches[38, 19]. These methods performed poorly when dealing with complex motions. As RNNs can naturally process sequential data, researchers widely used RNNs and their variants for motion prediction tasks[11, 9, 14, 31]. EDR[9] maps motion sequences to a latent space using a nonlinear encoder and models temporal features using LSTM (Long Short-Term Memory Network). To avoid ambiguity when predicting long-term motions, VAE-LSTM[14] introduces the encoding of control signals into a variational inference framework trained to learn human motion manifolds. United[31] achieves smooth prediction by modeling first-order motion derivatives. To incorporate spatial information modeling, Structure-RNN[18] unfolds the spatiotemporal graph in the temporal dimension and decomposes it into a set of contributing factor components fed into RNN modeling. SkelNet[13], at each time frame, obtains local representations of human structure through special component layer branches.
Feed-forward Methods
The step-by-step modeling and generation approach of RNNs often leads to the issue of error accumulation. In recent times, many methods have adopted a feed-forward paradigm, utilizing GNN and Transformer models, to generate prediction sequences in a non-autoregressive manner[26, 3, 10, 30, 34]. These approaches leverage graph structures and attention computations that adapt to the non-Euclidean distribution properties of joints in space and time. LTD[28], HisRep[29], SGSN[22], and SPGSN[24] transform motion sequences to the frequency domain, modeling the trajectories of joints. DMGNN[20], MST-GNN[23], MSR-GCN[7], Sybio-GNN[21], and MMotion-Att[30] learn motion patterns at multiple levels by modeling human representations at different scales. DGMGNN[20], MANet[34], DD-GCN[40], and DSTD-GCN[10] model the dynamic characteristics of joints, adapting to the dynamic properties of motion. Transformer-based models utilize self-attention computations, which enhance the model’s capability to capture long-range dependencies between joints[1, 32, 41]. Among them, AuxFormer[41] achieves better results by incorporating denoising and reconstruction tasks to assist the prediction task. In addition, lightweight models based on MLP[2, 12] have also demonstrated success.
@@ 2.2 Self-supervised Pretraining
Self-supervised pretraining has brought about a revolution in the understanding and generation of language and images. This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT[8], RoBERTa[25] and other models. For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method[15, 5]. MAE[15] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features. The self-learning paradigm of autoencoder in MAE is also widely used in multimodal[4], video processing[36], point cloud data processing[42] and other tasks.
@@ 3 Method
Figure 2: Detail of the model architecture. Our proposed model comprises Past Motion Encoder (PME) and Future Motion Predictor (FMP), where the former consists of $N\times$ SpatioTemporal Attention (ST-Attention), and the latter consists of $M$ ST-Attention and Past Motion Guided Attention (PMG-Attention). Here, the finetune stage is taken as an example to show the model structure (the pretrain stage only distinguishes the number of forward propagation of PME and incomplete input sequences).
@@ 3.1 Problem Background
The purpose of 3D skeleton-based human motion prediction is to generate the unseen and consecutive future motion based on past motion patterns. Assume a historical motion sequence consisting of $T$ frames in the Cartesian coordinate system, denoted as $\mathbf{X}^{P}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ represents the number of joints and $K$ is the feature dimension (the initial value of $K$ is 3, representing the $x,y,z$ coordinates). Assume $\mathbf{X}^{F}=[\textbf{x}^{1},\textbf{x}^{2},...,\textbf{x}^{L}]\in\mathbb{R}^{L\times J\times K}$ is the future motion pose of $L$ frames. Our goal is to design and train a network $\mathcal{P}(\cdot)$ that takes $\mathbf{X}^{P}$ as input and $\mathbf{X}^{F}$ as output, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P},\pi_{\theta})$, where $\pi_{\theta}$ is the trainable parameters.
@@ 3.2 Model Architecture
As shown in Fig. 2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Predictor (FMP). The PME aims to extract spatiotemporal features from the past motion sequence. The FMP aims to generate the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
@@
where $\mathbf{W}\in\mathbb{R}^{K\times C},\mathbf{b}\in\mathbb{R}^{C}$ are the weights and biases of the linear layer, and $\mathbf{P}_{t},\mathbf{P}_{s}$ are the temporal and spatial position encoding, which are randomly initialized and trainable. $C$ denotes the number of feature channels. For the output of PME and FMP, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
@@
We design PME to model the features of the past motion sequence. Human motion sequences contain two features: the change of joint position over time and the relationship between joints in space. We use the SpatioTemporal Attention (ST-Attention) to simultaneously model the spatiotemporal information of the joints. Specifically, PME contains $N$ cascaded ST-Attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension. Assuming the input to the $l^{th}$ layer of ST-Transformer is $\mathbf{H}^{l}$, the computation process can be expressed as:
@@
Future Motion Predictor (FMP)
The purpose of FMP is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features). FMP comprises two modules: the ST-Attention module for modeling the sequence dependencies within itself and the Past Movement Guided Attention module (PMG-Attention). PMG-Attention applies cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
@@
where $\mathbf{H}^{P}$ is the past motion features output by PME. $\mathbf{H}^{l}$ is the input of $l^{th}$ PMG-Attention. $\mathcal{F}_{c-attn}^{s}(\cdot)$ / $\mathcal{F}_{c-attn}^{t}(\cdot)$ represents the spatial/temporal multi-head cross-attention calculation with the bottle neck. We provide more information about the ST-Attention and PMG-Attention in the supplementary materials (Sec. 7).
@@ 3.3 Past Movement Guided Pretraining
We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the left part of Fig. 1(d)), which compromise three steps: 1) past motion reconstruction; 2) past motion encoding; 3) future motion reconstruction. We will introduce each step in the following.
@@
Given the past motion sequence $\mathbf{X}^{P}\in\mathcal{R}^{T\times J\times K}$, we first calculate the motion velocity of each joint $\mathbf{V}=\mathbf{X}_{2:T}^{P}-\mathbf{X}_{1:T-1}^{P}\in\mathcal{R}^{(T-1)\times J\times K}$. We find the threshold $s$ in $\mathbf{V}$ such that the $r$of the values in $\mathbf{V}$ are greater than the threshold $s$. The mask of the past motion sequence $\mathbf{M}^{P}$ can be mathematically expressed as:
@@
The masked past motion sequence can be computed by $\mathbf{X}_{masked}^{P}=\mathbf{X}^{P}\times\mathbf{M}^{P}$. Then, the joint embedders are applied to obtain the masked motion embeds $\mathbf{X}_{emb}^{P}$. Noting that we incorporate the trainable temporal and spatial position encoding into the motion embeds to learn the temporal order and spatial relationships of joints. Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Position Motion Encoder.
Past Motion Encoding
This step aims to compute the features of complete past motion, which will be used to guide the future motion reconstruction and integrate the hidden relationship between the past sequence and the future sequence into the network. Given the complete input sequence $\mathbf{X}^{P}$, we first apply joint embedder to obtain the motion embeds $\mathbf{X}_{emb}^{P}$. Then, the past motion features can be computed by $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$.
@@
Given the future motion sequence $\mathbf{X}^{F}\in\mathcal{R}^{L\times J\times K}$, we first use the same method for masking and embedding the future motion sequence as for processing the past motion sequence. The masked future motion embeds, denoted as $\mathbf{X}_{emb}^{F}$ and the complete past motion features $\mathbf{H}^{P}$ are the input of FMP. In the PMG-Attention of FMP, past motion features guide the reconstruction of masked future motion through cross-attention mechanism. Finally, the reconstructed motion sequence can be expressed as: $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMP}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P})))$.
3.4 Fine-tuning for Motion Prediction
In the finetune phase, we use the complete past motion sequence as input to the PME to get past motion features (as shown in the right part of Fig. 1(d)). In FMP, we use the initial state initialized with the value of 0 (but retain the position encoding) as input, and output the complete future sequence guided by the past motion features. Given the past motion sequence $\mathbf{X}^{P}$, this process can be represented as:

### v2

# MoReFun: Past-Movement Guided Motion Representation Learning for Future Motion Prediction and Understanding

@@ Abstract
3D human motion prediction aims to generate coherent future motions from observed sequences, yet existing end-to-end regression frameworks often fail to capture complex dynamics and tend to produce temporally inconsistent or static predictions—a limitation rooted in representation shortcutting, where models rely on superficial cues rather than learning meaningful motion structure. We propose a two-stage self-supervised framework that decouples representation learning from prediction. In the pretraining stage, the model performs unified past–future self-reconstruction, reconstructing the past sequence while recovering masked joints in the future sequence under full historical guidance. A velocity-based masking strategy selects highly dynamic joints, forcing the model to focus on informative motion components and internalize the statistical dependencies between past and future states without regression interference. In the fine-tuning stage, the pretrained model predicts the entire future sequence, now treated as fully masked, and is further equipped with a lightweight future-text prediction head for joint optimization of low-level motion prediction and high-level motion understanding. Experiments on Human3.6M, 3DPW, and AMASS show that our method reduces average prediction errors by 8.8% over state-of-the-art methods while achieving competitive future-motion understanding performance compared to LLM-based models. Code is available at: https://github.com/JunyuShi02/MoReFun
@@ 1 Introduction
Figure 1: Comparison of complex motion prediction results. While previous methods struggle with complex movements, producing poor and unnatural static poses, our method demonstrates improved accuracy by addressing the inherent weakness in representation capability, yielding a natural motion prediction.
3D human pose prediction, which aims to generate continuous and natural future motions based on historical observed sequences, is a pivotal technology for applications such as human-computer interaction and autonomous driving. Despite considerable strides made by recent studies, particularly those based on GNNs [ma2022progressively, cai2020learning, fu2023learning, mao2021multi, shi2023multi] and Transformers [aksan2021spatio, guo2023back, wei2025alien], existing state-of-the-art (SOTA) models exhibit fundamental failures in accurately predicting complex motions. Their predicted motions often rapidly degenerate into unnatural static poses or lose temporal coherence, as shown in the Fig. 1.
This failure is not attributable to specific models (e.g. Transformers) but rather stems from a fundamental flaw in the prevailing end-to-end regression paradigm used for motion prediction. Current methods frame this task as a direct regression problem. This approach compels the model to simultaneously learn complex spatio-temporal representations and regress a highly uncertain future within a single training stage. This results in a phenomenon we term “Representation Shortcutting” where the model learns only superficial features sufficient to minimize the regression loss on the training data, while failing to capture the underlying physical principles and joint constraints governing motion.
To improve the prediction performance, researchers have pursued two main paths. The first path involves refining the regression paradigm. For instance, feed-forward architectures [ma2022progressively, xu2023eqmotion, sun2024moml] have been used to mitigate temporal error accumulation, while other methods model low-frequency DCT signals to reduce motion jitter [wang2024gcnext, li2021skeleton]. Moreover, the system latency [sun2025lal] and unexpected perturbation [yue2024human] are also under consideration. However, these methods remain constrained by the “Representation Shortcutting” problem. The second path explores auxiliary learning, such as using parallel denoising tasks to aid prediction [xu2023auxiliary]. Yet, this parallel multi-task paradigm has critical flaws: (1) The auxiliary and main (motion prediction) tasks may suffer from mutual interference; and (2) It fails to explicitly model the fundamental dependency of future motion on past observations, which is central to accurate motion prediction.
In this work, we decouple representation learning from the prediction task into two distinct stages. In the first stage, we propose a self-supervised learning (SSL) framework based on past-guided future reconstruction to pretrain the motion model. We mask joints within the future motion sequence and compel the model to leverage the complete past sequence as guidance to reconstruct the masked elements. This SSL approach enables the model to acquire rich motion priors, free from the interference of a regression loss, helping it grasp the transition logic and statistical dependencies between past and future states. Concurrently, we also perform self-reconstruction on the past sequence to enhance the encoder’s representational power.
To maximize the efficacy of this SSL stage, we introduce a velocity-based masking strategy. Unlike random masking, we prioritize masking joints that exhibit the most significant movement between adjacent frames. This forces the model to focus its representational capacity on the most informative and dynamic components of the motion (e.g., the upper limbs during an “eating” action), rather than on low-information, static joints.
In the second stage, we fine-tune the pretrained model on the downstream motion prediction task. Here, we treat the entire future sequence as “masked” and task the model with its recovery. Furthermore, we explore the synergistic relationship between motion prediction and motion understanding. We design a novel joint fine-tuning objective where the model, given the history, is tasked with simultaneously predicting: (1) the future motion sequence (low-level prediction) and (2) a future motion description (high-level understanding). We find that these two tasks are mutually beneficial: the high-level, textual understanding of the future movement constrains and refines the low-level motion prediction, and vice versa.
Our key contributions can be summarized as follows:
  • We propose a self-supervised motion representation learning framework that unifies past and future self-reconstruction, with future reconstruction guided by historical motion cues. A velocity-based masking strategy is applied to emphasize dynamic joints during reconstruction. This paradigm enables the model to capture intrinsic spatiotemporal structure as well as the dependencies between past and future states.
  • We develop a joint fine-tuning objective that tasks the model with predicting both the future motion sequence and future textual description. This coupling of low-level motion prediction and high-level semantic understanding yields mutually reinforcing improvements in coherence and realism.
@@
  • We conduct extensive experiments and demonstrate that our method reduces the average prediction errors by 8.8% compared to state-of-the-art methods on the 3DPW, Human3.6M, and AMASS datasets. Moreover, with a lightweight text prediction head, our model achieves competitive performance in future motion understanding compared to LLM-based methods [wu2025mg].
@@ 2 Related Work
Figure 2: Overall architecture of MoReFun, which consists of a Past Motion Encoder (PME) and a Future Motion Decoder (FMD). The PME is built with stacked Spatial and Temporal Self-Attention layers, processing the past motion sequence. The FMD receives masked future motion together with the encoded past-motion features, and is composed of Spatial/Temporal Cross-Attention followed by Spatial/Temporal Self-Attention. Through these stages, the decoder leverages the past-motion representation to guide the reconstruction of masked future joints.
@@
Early motion prediction methods relied on probabilistic models [wang2007gaussian, lehrmann2014efficient], yet their strong assumptions limited their ability to capture complex human dynamics. With deep learning, RNN-based approaches [ghosh2017learning, fragkiadaki2015recurrent, habibie2017recurrent, martinez2017human] became dominant by encoding historical trajectories and autoregressively generating future poses, though they often suffer from error accumulation during long-term prediction. To overcome this, recent works shift toward feed-forward architectures that directly predict future sequences in a non-autoregressive manner. Graph-based models [ma2022progressively, li2021multiscale, dang2021msr, li2021symbiotic, shi2023multi] leverage skeletal topology, frequency-domain representations [mao2019learning, mao2020history, shi2024gradient] provide compact trajectory modeling, and dynamic GNNs [li2020dynamic, fu2023learning] adapt joint relations to motion changes. Transformer-based models [aksan2021spatio, martinez2021pose, xu2023auxiliary] further enhance long-range dependency modeling, with AuxFormer [xu2023auxiliary] improving prediction using auxiliary denoising and reconstruction tasks. Meanwhile, lightweight MLP designs [bouazizi2022motionmixer, guo2023back] demonstrate that strong performance can also be achieved without heavy attention mechanisms. Together, these feed-forward approaches alleviate autoregressive drift and have become the mainstream paradigm for modern human motion prediction.
@@ 2.2 Self-supervised Pretraining
Self-supervised pretraining has brought about a revolution in the understanding and generation of language and images. This technique serves as one of the fundamental training methods of Large Language Models (LLMs), which is widely used in BERT [devlin2018bert], RoBERTa [liu2019roberta] and other models. For image processing tasks, recovering masked pixels/tokens is a common self-supervised pretraining method [he2022masked, chen2024context]. MAE [he2022masked] proves that even when a substantial portion of image information is masked, the network can reconstruct a complete image based on a limited number of visible features. The self-learning paradigm of autoencoder in MAE is also widely used in multimodal [chen2023pimae], video processing [tong2022videomae], motion generation [shi2025genm], point cloud data processing [zhang2023learning] and other tasks.
2.3 Motion Captioning
Motion captioning aims to generate natural language descriptions for 3D human motion sequences [petrovich2022temos, shi2025mogic]. Recent LLM-based approaches [li2025lamp, jiang2023motiongpt] further leverage powerful pretrained language models to produce more accurate and semantically coherent motion descriptions. Despite these advances, the role of future-motion semantics in enhancing motion prediction remains largely underexplored, leaving a gap between motion understanding and predicting.
@@ 3.1 Problem Background
3D skeleton-based human motion prediction aims to generate unseen and consecutive future motions from historical observations. Given a past motion sequence $\mathbf{X}^{P}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times J\times K}$, where $J$ denotes the number of joints and $K$ is the feature dimension (typically $K=3$ for the $x,y,z$ coordinates), the goal is to predict the future motion sequence $\mathbf{X}^{F}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}]\in\mathbb{R}^{L\times J\times K}$. We therefore design a predictive model $\mathcal{P}(\cdot)$ that maps the observed past motion to its future continuation, i.e., $\mathbf{X}^{F}=\mathcal{P}(\mathbf{X}^{P};\pi_{\theta})$ with trainable parameters $\pi_{\theta}$. Beyond pose prediction, we further consider the generation of a semantic description of the future behavior. Given the same historical sequence, the model outputs a future text sequence $\mathbf{Y}^{F}=[y_{1},\ldots,y_{D}]$ that summarizes the expected motion. This formulation naturally extends the task to a joint prediction problem, where the model produces both future motion and its corresponding semantic description: $(\mathbf{X}^{F},\mathbf{Y}^{F})=\mathcal{P}(\mathbf{X}^{P};\pi_{\theta})$.
Figure 3: Overview of the finetuning stage. The Past Motion Encoder extracts spatiotemporal representations from the observed history motion. The shared past-motion feature is fed into two branches: (1) the Future Motion Decoder, which predicts the future 3D pose sequence from a zero-initialized state, and (2) the Understanding Decoder, which produces a natural-language description of the upcoming motion. Both branches are jointly optimized with motion and text supervision.
@@ 3.2 Model Architecture
As shown in Fig. 2, our proposed model consists of a Past Motion Encoder (PME) and a Future Motion Decoder (FMD). The PME aims to extract spatiotemporal features from the past motion sequence. The FMD aims to predict the future motion sequence based on the initial state (which consists of zero-valued inputs in motion prediction tasks) and past motion features.
@@
where $\mathbf{W}\in\mathbb{R}^{K\times C},\mathbf{b}\in\mathbb{R}^{C}$ are the weights and biases of the linear layer, and $\mathbf{P}_{t},\mathbf{P}_{s}$ are the temporal and spatial position encoding, which are randomly initialized and trainable. $C$ denotes the number of feature channels. For the output of PME and FMD, a linear layer as the Predict Head is applied to obtain the 3D motion sequence.
@@
We design PME to model the features of the past motion sequence. Human motion sequences contain two features: the change of joint position over time and the relationship between joints in space. We use the spatiotemporal attention to simultaneously model the spatiotemporal information of the joints. Specifically, PME contains $N$ cascaded spatiotemporal attention modules, where each module first models the sequence dependencies in the temporal dimension and then models the joint relationships in the spatial dimension. Assuming the input to the $l^{th}$ layer is $\mathbf{H}^{l}$, the computation process can be expressed as:
@@
Future Motion Decoder (FMD)
The purpose of FMD is to generate future motion sequences based on the initial state and the output from PME (i.e., past motion features). FMD comprises two modules: the spatiotemporal attention module for modeling the sequence dependencies within itself and the spatiotemporal cross-attention to incorporate past motion features into the current state, which can be mathematically represented as:
@@
where $\mathbf{H}^{P}$ is the past motion features output by PME. $\mathbf{H}^{l}$ is the input of $l^{th}$ layer of FMD. $\mathcal{F}_{c-attn}^{s}(\cdot)$ / $\mathcal{F}_{c-attn}^{t}(\cdot)$ represents the spatial/temporal multi-head cross-attention calculation with the bottle neck.
@@ 3.3 Past Movement Guided Pretraining
We apply the self-reconstruction pretraining paradigm to help the model learn the spatiotemporal representation of joints in motion sequences (as shown in the Fig. 2), which compromise two proxy tasks: 1) past motion reconstruction; 2) future motion reconstruction. We will introduce each step in the following.
@@
Given the past motion sequence $\mathbf{X}^{P}\in\mathcal{R}^{T\times J\times K}$, we first calculate the motion velocity of each joint $\mathbf{V}=\mathbf{X}_{2:T}^{P}-\mathbf{X}_{1:T-1}^{P}\in\mathcal{R}^{(T-1)\times J\times K}$. We find the threshold $s$ in $\mathbf{V}$ such that the rr% of the values in $\mathbf{V}$ are greater than the threshold $s$. The mask of the past motion sequence $\mathbf{M}^{P}$ can be mathematically expressed as:
@@
The masked past motion sequence can be computed by $\mathbf{X}_{masked}^{P}=\mathbf{X}^{P}\times\mathbf{M}^{P}$. Then, the joint embedders are applied to obtain the masked motion embeds $\mathbf{X}_{emb}^{P}$. Finally, we calculate the reconstructed motion sequence $\mathbf{X}_{rec}^{P}=\mathcal{F}_{PH}(\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P}))$, where $\mathcal{F}_{PH}(\cdot)$ and $\mathcal{F}_{PME}(\cdot)$ are Predict Head and Past Motion Encoder.
Through this past-motion self-reconstruction objective, the model learns to capture both the spatial structure within individual poses and the temporal dependencies across consecutive motions. This encourages the encoder to acquire rich and robust 3D human motion representations that benefit downstream tasks.
@@
Given the complete past motion sequence $\mathbf{X}^{P}$, we first apply the joint embedder to obtain the motion embeddings $\mathbf{X}_{emb}^{P}$, which are then encoded into past-motion features via $\mathbf{H}^{P}=\mathcal{F}_{PME}(\mathbf{X}_{emb}^{P})$. For the future motion sequence $\mathbf{X}^{F}\in\mathbb{R}^{L\times J\times K}$, we adopt the same velocity-based masking strategy and embedding procedure, producing the masked embeddings $\mathbf{X}_{emb}^{F}$. The masked future-motion embeddings $\mathbf{X}_{emb}^{F}$ and the past-motion features $\mathbf{H}^{P}$ are fed into the FMD module, where the PMG-Attention uses cross-attention to guide the reconstruction of masked future motion. The final reconstructed future motion is given by $\mathbf{X}_{rec}^{F}=\mathcal{F}_{PH}(\mathcal{F}_{FMD}(\mathbf{X}_{emb}^{F},\mathbf{H}^{P}))$.
In this reconstruction task, the model completes masked future motions based on past observations. This encourages it to explicitly learn the fundamental dependency of future motion on past inputs, as well as the underlying priors of human movement patterns and motion trajectories, facilitating downstream tasks.
3.4 Fine-tuning for Prediction and Understanding
In the finetuning phase, we use the complete past motion sequence as input to the PME to obtain the past motion features (as shown in the Fig.3). In the FMD, we initialize the future query tokens with zeros (while preserving positional encodings) and decode the full future motion sequence guided by the past features. Given the past motion sequence $\mathbf{X}^{P}$, the prediction process is formulated as: